# LLMs for Generating and Evaluating Counterfactuals: A Comprehensive Study

### v1

@@ 4.3 LLMs for CFs Evaluation
We examine how reliable are LLMs for CFs evaluation by asking them to evaluate two sets of human CFs: an honest set and a corrupted set. The “honest set” refers to a collection of human CFs, for which the ground truth labels are provided, whereas the “corrupted set” consists of instances, for which wrong labels differing from the gold labels are provided. In the context of NLI, the third label, distinct from both the target and factual labels, is selected for inclusion in the corrupted set. For SA, the reverse label is chosen while the factual label remains undisclosed. Initially, we prompt GPT3.5 and GPT4 to assess whether the provided CFs accurately represent the target labels by assigning a score from 1 to 4 (cf. Appendix F). Here, a score of 1 or 2 indicates disagreement (complete or partial) with the target label, while a score of 3 or 4 indicates agreement (partial or complete) with the target label. Ideally, the evaluation LLMs should give high scores to the honest set, and low scores to the corrupted set. We show the distributions for disagreements and agreements in Table 6.
On SA, both LLMs perform well, but GPT4 exhibits higher sensitivity to the corrupted examples. On NLI, we notice that GPT3.5 gives high flip label scores to humans CFs with both correct and incorrect labels. GPT4 performs much better, but still exhibits high tendency to agree with wrong labels ($\sim$ 40%). The results can be explained by the tendency of LLMs to agree with the provided answers, especially on reasoning tasks Zheng et al. (2023). To verify this, we prompt both LLMs to classify the same set of NLI CFs by choosing one the three labels (entailment, neutral, contradiction) using a similar prompt. The classification results in Table 7 show an improved performance compared to asking the same LLMs if they agree with incorrect labels (cf. Table 6). We also compare the flip label score distributions of GPT3.5 and GPT4 on the corrupted set in Table 10, and observe that even though GPT3.5 gives high scores to corrupted inputs it is less certain (most frequent score is 3), whereas GPT4 tends to be more certain and assigns mostly 1 or 4 (> 93%).
@@
We conduct a wide-scale CFs evaluation with GPT4. Besides verifying the target label FL, we also ask GPT4 to judge if there are any unnecessary alterations UA, and if the CF is realistic RS. For these aspects, we use a scoring scheme ranging from 1 to 4, where higher scores indicate better performance. The results for GPT-4 evaluation can be found in Table 8.
The evaluation scores from GPT4 show that GPT4 CFs are the best overall. On SA, Mistral 7b comes in second place after GPT4, whereas on NLI human CFs take the second position. GPT4 might have a bias to prefer its own generations Panickssery et al. (2024). The scores for SA are lower, which might be explained by the fact that SA is an easier task than NLI, and therefore evaluating SA CFs is also easier for GPT4, and potentially more accurate. To further verify the evaluation scores from GPT4, we calculate the correlations between GPT4 scores and the scores from the intrinsic evaluation. The correlations shown in Table 9 reflect strong relation on SA with respect to all three aspects (flipping label, minimal changes, realisticness). On NLI, we notice the same pattern except for realisticness/perplexity, where we notice a weak negative relation. This might be due to the nature of NLI instances that are short image captions, and are not necessarily grammatical.
LLMs show a high tendency to agree with the provided labels even if these are incorrect, especially on tasks that require reasoning such as NLI. GPT4 evaluation scores correlate with automatic metrics for flipping label, distance and fluency.
LLM/Set Task 1/2 3/4 Avg.
@@
LLM FL UA RS Avg.
SA
Human Crowd 3.66 2.95 2.58 3.06
Human Expert 3.54 2.69 2.49 2.91
GPT3.5 3.58 2.91 2.65 3.05
GPT4 3.79 3.15 2.91 3.28
LLAMA2 7B 3.60 2.74 2.63 2.99
LLAMA2 70B 3.70 2.75 2.47 2.97
Mistral 7B 3.85 2.84 2.69 3.13
Mistral 56B 2.58 1.74 1.75 2.02
hypothesis
Human Crowd 3.81 3.96 3.81 3.86
GPT3.5 3.19 3.93 3.74 3.62
GPT4 3.96 3.98 3.92 3.95
LLAMA2 7B 3.23 3.74 3.66 3.54
LLAMA2 70B 3.49 3.68 3.56 3.58
Mistral 7B 3.50 3.70 3.65 3.62
Mistral 56B 3.02 3.45 3.48 3.32
premise
Human Crowd 3.58 3.88 3.86 3.77
GPT3.5 2.51 3.82 3.69 3.34
GPT4 3.68 3.83 3.84 3.78
LLAMA2 7B 2.96 3.38 3.67 3.34
LLAMA2 70B 3.35 3.46 3.67 3.49
Mistral 7B 2.97 3.63 3.74 3.45
Mistral 56B 2.37 3.11 3.49 2.99
@@ 4.4.1 CFs for NLI
  • Copy-Paste: When asked to generate a CF, and change the label form contradiction to entailment, LLMs will use the unchanged part (premise or hypothesis) as output. This a clever but lazy way to flip the label to entailment, since two identical sentences would naturally have the label entailment. These CFs were given perfect scores by GPT4. Table 12 in the Appendix shows the percentage of copy-paste CFs for all LLMs (at most 4.27% for GPT3.5).
  • Negation: When asked to to change the label from entailment to contradiction, LLMs would attempt to negate the premise/hypothesis. The negations does not make sense in the observed CFs, and GPT4 rightly assigns them the lowest scores possible.
  • Inconsistency: These examples contain contradictory or illogical sentences. Here, GPT4 sometimes incorrectly assigned these CFs high scores.
We show some examples alongside the assigned categories in Table 13 in the Appendix.
@@ 5 Related Work
There exists several methods for generating CFs, but most of them are desigend for a specific LLM. The CFs generated by MICE Ross et al. (2021) are intended for debugging models, and not for data augmentation. Polyjuice Wu et al. (2021) requires specifying the type of edits that should be conducted, and the resulting CFs should be manually labeled. Robeer et al. (2021). DISCO Chen et al. (2023) uses GPT-3’s fill-in-the-blanks mode, which is unavailable in most open source LLMs and would require adapting them. CREST Treviso et al. (2023) depends on a rationalizer module and the editor module is a masked LM that needs to be further trained. Instead, we decided to prompt LLMs to generate CFs by providing instructions and an example. We provide more details in Section 3.2.
@@ 6 Conclusion
In this work, we investigated the use of various LLMs for CFs generation. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for SA is less challenging than NLI, where LLMs show weaknesses in generating CFs that change the original label. CFs from LLMs can replace human CFs for the purpose of data augmentation on SA and achieve similar performance, while on NLI further improvements are needed. Furthermore, our results suggest that CFs with minimal changes are essential for data augmentation. We also showed that when asked to asses CFs, LLMs exhibit a strong bias towards agreeing with the provided label even if this label is incorrect. GPT4 appears to be more robust than GPT3.5 against this bias, and its scores correlate well with automatic metrics. However, some failures still exist as shown in our qualitative analysis, indicating limitations of LLMs that warrant further investigation. Future work should focus on leveraging LLMs for higher quality NLI CFs, which correctly changes the label and keep minimal changes, and assessing the evaluation abilities of LLMs in mislabeled data settings.
@@ 7 Limitations
Appendix A Robustness Evaluation
CFs are used to construct Constrast Sets, which are leveraged to conduct fine-grained evaluations, allowing the detection of model failures Gardner et al. (2020). Similarly, we construct Constrast Sets with different LLMs and evaluate how task-specific models perform on these sets, comparing them to human-generated Contrast Sets. High quality CFs should be challenging for classifiers in order to help users detect failure modes in their models.
Test split Success Rate Acc.
@@ Appendix A Robustness Evaluation
orig NA 90.16
crowd NA 92.21
experts NA 86.27
gpt3.5 100.0% 90.37
gpt4 99.59% 96.09
llama2 7b 98.98% 86.34
llama2 70b 66.6% 96.31
mistral 7b 84.22% 97.32
mistral 56b 56.35% 68.73
@@
crowd NA 39.88
gpt3.5 100.0% 31.00
gpt4 100.0% 41.88
llama2 7b 96.0% 31.90
llama2 70b 15.25% 34.43
mistral 7b 96.12% 26.92
mistral 56b 64.75% 24.71
@@
crowd NA 58.50
gpt3.5 100.0% 32.50
gpt4 100.0% 47.50
llama2 7b 99.62% 34.38
llama2 70b 21.75% 37.36
mistral 7b 94.38% 31.92
mistral 56b 64.5% 37.40
Table 11: Robustness Evaluation on CFs from different LLMs. Classification model is trained only on original data from Kaushik et al. (2020). Evaluation is done CFs from crowd and expert humans, and from various LLMs.
A.1 Robustness Evaluation
We train and optimize a BERT-base model on the original training data from Kaushik et al. (2020), and compare its performance on CFs from crowd and expert humans, and different LLMs. For details on hyperparameter optimization, we refer to Appendix C. Higher quality CFs should be more challenging for the model, and result in a performance drop. In addition to accuracy, we report the success rate in generating CFs. If the output of the LLM does not correspond to the template we provide in the demonstration example it is discarded and considered a failed generation. The results for both, SA and NLI, are shown in Table 11.
In general, we notice that GPT models are the most stable, as they are able to generate CFs in most cases. Counter-intuitively, the smaller variants of mistral and llama2 are more successful in generating CFs than their larger counterparts. One reason for this could be that larger models tend to be undertrained. Even though the model size and the amount of training data should be scaled equally Hoffmann et al. (2022), llama2 7b and llama2 70b are pre-trained on the same amount of data.  [^4] See Table 1 in Touvron et al. (2023)
On SA, we notice that the crowd CFs are easier to classify than the original test set for our optimized classifier (accuracy increases 90.16 $\rightarrow$ 92.21). Conversely, the performance drops on expert CFs (90.16 $\rightarrow$ 86.27). This confirms the high quality of expert CFs. The CFs from LLMs are not challenging (accuracy >= 90) with two exceptions: llama2 7b and mistral 56b. One cause for this might be that these two LLMs are generating non-sensible CFs.
On NLI, the CFs are generated twice: once for the premise part and once for the hypothesis part. In each case, the other (factual) part is kept. The crowd CFs are more difficult for the model than the original test set, especially when the premise is changed (73.50 $\rightarrow$ 39.88 (premise), 58.50 (hypothesis)). The LLMs-generated CFs appear to be even more challenging (accuracy drops to 24.71 at worst). LLMs might be able to generate diverse CFs, which are not typically generated by humans Wu et al. (2021), and are thus challenging for the classifier. Another reason could be that CFs from LLMs have low quality.
A major limitation of this experiment is the abscence of ground truth labels.
@@ Appendix B Further Analysis
gpt4 premise 4.14
llama2 hypothesis 2.26
llama2 premise 3.00
llama2_70b hypothesis 2.87
llama2_70b premise 1.64
mistral hypothesis 1.59
mistral premise 0.91
mistral_56b hypothesis 2.72
mistral_56b premise 1.17
@@ Appendix D Examples
original_sentence1 original_sentence2 contrast text original label target label FL UA RS LLM part
Copy-Paste
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table on the floor in a building that overlooks the city. contradiction entailment 4.0 4.0 4.0 mistral hypothesis
Inconsistency
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 mistral premise
An old man in a blue shirt and white pants glowers by a glass door. A man resting outside A glass man by a door. neutral contradiction 1.0 1.0 1.0 crowd hypothesis
Negation
A solo shot of a man riding a four wheeler in an atv event. The man is entered in a event. A solo shot of a man rode a four wheeler in an no atv event. The man is not entered in a no event. entailment contradiction 1.0 1.0 1.0 mistral_56b premise
A boy dressed for summer in a green shirt and kahki shorts extends food to a reindeer in a petting zoo. A boy alien dressed for summer in a green shirt and kahki shorts A girl dressed for winter in a red shirt and black shorts withdraws water from a shark in an aquarium entailment contradiction 1.0 1.0 1.0 llama2_70b premise
A man with a beard is talking on the cellphone and standing next to someone who is lying down on the street. A man is prone on the street while another man stands next to him. Unproning on the street while another man lies in bed. entailment contradiction 1.0 1.0 1.0 mistral_56b hypothesis
Table 13: Categorization of a sample of incorrect NLI CFs.
original_sentence1 original_sentence2 contrast text original label target label FL UA RS LLM part
@@
A photographer and assistant take a photo of a man in a suit and a woman in a dress in front of large white columns and a yellow wall. The photographer is taking a photo of a dog. The photographer takes a photo of a dog and a man in a suit and a woman in a dress contradiction entailment 4.0 4.0 4.0 llama2_70b revised_hypothesis
@@
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 mistral revised_premise
A boy dressed for summer in a green shirt and kahki shorts extends food to a reindeer in a petting zoo. A boy alien dressed for summer in a green shirt and kahki shorts A girl alien dressed for summer in a green shirt and kahki shorts entailment contradiction 1.0 1.0 1.0 llama2_70b revised_hypothesis
@@
Appendix E Data
Some statistics about the number of generated CFs are shown in Table 15.
task llm split size
nli NLI_revised_hypothesis dev 400
nli NLI_revised_hypothesis test 800
nli NLI_revised_hypothesis train 3332
nli NLI_revised_premise dev 400
nli NLI_revised_premise test 800
nli NLI_revised_premise train 3332
nli gpt3.5_revised_hypothesis dev 400
nli gpt3.5_revised_hypothesis test 800
nli gpt3.5_revised_hypothesis train 3332
nli gpt3.5_revised_premise dev 400
nli gpt3.5_revised_premise test 800
nli gpt3.5_revised_premise train 3330
nli gpt4_revised_hypothesis dev 400
nli gpt4_revised_hypothesis test 800
nli gpt4_revised_hypothesis train 3329
nli gpt4_revised_premise dev 400
nli gpt4_revised_premise test 800
nli gpt4_revised_premise train 3332
nli llama2-7b_revised_hypothesis dev 335
nli llama2-7b_revised_hypothesis test 797
nli llama2-7b_revised_hypothesis train 3312
nli llama2-7b_revised_premise dev 280
nli llama2-7b_revised_premise test 768
nli llama2-7b_revised_premise train 3198
nli llama2_70b_revised_hypothesis dev 116
nli llama2_70b_revised_hypothesis test 174
nli llama2_70b_revised_hypothesis train 582
nli llama2_70b_revised_premise dev 84
nli llama2_70b_revised_premise test 122
nli llama2_70b_revised_premise train 481
nli mistral-7b_revised_hypothesis dev 380
nli mistral-7b_revised_hypothesis test 755
nli mistral-7b_revised_hypothesis train 3150
nli mistral-7b_revised_premise dev 379
nli mistral-7b_revised_premise test 769
nli mistral-7b_revised_premise train 3174
nli mistral_56b_revised_hypothesis dev 257
nli mistral_56b_revised_hypothesis test 516
nli mistral_56b_revised_hypothesis train 2076
nli mistral_56b_revised_premise dev 270
nli mistral_56b_revised_premise test 518
nli mistral_56b_revised_premise train 2126
sentiment experts dev 245
sentiment experts test 488
sentiment crowd dev 245
sentiment crowd test 488
sentiment crowd train 1707
sentiment gpt3.5 new dev 245
sentiment gpt3.5 new test 488
sentiment gpt3.5 new train 1707
sentiment gpt4 new dev 245
sentiment gpt4 new test 486
sentiment gpt4 new train 1704
sentiment llama2-7b new dev 243
sentiment llama2-7b new test 483
sentiment llama2-7b new train 1681
sentiment llama2_70b new dev 88
sentiment llama2_70b new test 325
sentiment llama2_70b new train 738
sentiment mistral-7b new dev 200
sentiment mistral-7b new test 411
sentiment mistral-7b new train 1439
sentiment mistral_56b new dev 150
sentiment mistral_56b new test 275
sentiment mistral_56b new train 1026
Table 15: Statistics of datasets used and generated
Appendix F Prompt

### v2

@@ 4.3 LLMs for CFs Evaluation
We examine how reliable are LLMs for CFs evaluation by asking them to evaluate two sets of human CFs: an honest set and a corrupted set. The “honest set” refers to a collection of human CFs, for which the ground truth labels are provided, whereas the “corrupted set” consists of instances, for which wrong labels differing from the gold labels are provided. In the context of NLI, the third label, distinct from both the target and factual labels, is selected for inclusion in the corrupted set. For SA, the reverse label is chosen while the factual label remains undisclosed. Initially, we prompt GPT3.5 and GPT4 to assess whether the provided CFs accurately represent the target labels by assigning a score from 1 to 4 (cf. Appendix E). Here, a score of 1 or 2 indicates disagreement (complete or partial) with the target label, while a score of 3 or 4 indicates agreement (partial or complete) with the target label. Ideally, the evaluation LLMs should give high scores to the honest set, and low scores to the corrupted set. We show the distributions for disagreements and agreements in Table 7.
On SA, both LLMs perform well, but GPT4 exhibits higher sensitivity to the corrupted examples. On NLI, we notice that GPT3.5 gives high flip label scores to humans CFs with both correct and incorrect labels. GPT4 performs much better, but still exhibits high tendency to agree with wrong labels ($\sim$ 40%). The results can be explained by the tendency of LLMs to agree with the provided answers, especially on reasoning tasks Zheng et al. (2023). To verify this, we prompt both LLMs to classify the same set of NLI CFs by choosing one of the three labels (entailment, neutral, contradiction) using a similar prompt. The classification results in Table 8 show an improved performance compared to asking the same LLMs if they agree with incorrect labels (cf. Table 7). We also compare the flip label score distributions of GPT3.5 and GPT4 on the corrupted set in Table 11, and observe that even though GPT3.5 gives high scores to corrupted inputs it is less certain (most frequent score is 3), whereas GPT4 tends to be more certain and assigns mostly 1 or 4 (> 93%).
@@
We conduct a wide-scale CFs evaluation with GPT4. Besides verifying the target label FL, we also ask GPT4 to judge if there are any unnecessary alterations UA, and if the CF is realistic RS. For these aspects, we use a scoring scheme ranging from 1 to 4, where higher scores indicate better performance. The results for the GPT4 evaluation can be found in Table 9.
The evaluation scores from GPT4 show that GPT4 prefers LLM-CFs, and especially its own generations, which are given the highest scores on most datasets. On SA, Mistral 56B scores the highest with LLAMA 70B and GPT4 having slightly lower scores. On NLI, human CFs take the second position after GPT4. On HS, GPT4 performs the best, while human CFs are given the second lowest score on average. GPT4 might have a bias to prefer its own generations Panickssery et al. (2024). We further investigate this bias in Section 4.4.3. To further verify the evaluation scores from GPT4, we calculate the correlations between GPT4 scores and the scores from the intrinsic evaluation.
The correlations in Table 10 indicate strong correlation for label flipping on SA and NLI, but weak correlation on HS. This suggests that GPT4 highly agrees with the classifier. Minimal changes show weak correlation with Levenshtein distance on SA and HS, with moderate correlation on NLI, implying that GPT4 is not necessarily sensitive to small changes. GPT4 shows weak to moderate positive correlation on realisticness with perplexity on HS and SA, and moderate negative correlation on NLI. This discrepancy might be due to the nature of the different texts, i.e., while SA contains long movie reviews, NLI contains short image captions and HS contains tweets.
LLMs show a high tendency to agree with the provided labels even if these are incorrect, especially on tasks that require reasoning such as NLI. The correlation between GPT4 evaluation scores and automated metrics for label flipping, textual distance, and fluency varies across tasks .
LLM/Set Task 1&2 3&4 Avg.
@@
SA NLI - premise NLI - hypothesis Hate Speech
FL UA RS Avg. FL UA RS Avg. FL UA RS Avg. FL UA RS Avg.
Expert Humans 3.54 2.69 2.49 2.91 - - - - - - - - - - - -
Crowd Humans 3.66 2.95 2.58 3.06 3.58 3.88 3.86 3.77 3.81 3.96 3.81 3.86 3.04 3.54 3.19 3.26
GPT3.5 3.58 2.91 2.65 3.05 2.51 3.82 3.69 3.34 3.19 3.93 3.74 3.62 1.78 3.58 3.02 2.79
GPT4 3.79 3.15 2.91 3.28 3.68 3.83 3.84 3.78 3.96 3.98 3.92 3.95 3.65 3.73 3.63 3.67
LLAMA2 7B 3.60 2.74 2.63 2.99 2.96 3.38 3.67 3.34 3.23 3.74 3.66 3.54 - - - -
LLAMA2 70B 3.87 3.05 2.96 3.29 3.07 3.68 3.77 3.51 3.60 3.89 3.75 3.75 - - - -
Mistral 7B 3.85 2.84 2.69 3.13 2.97 3.63 3.74 3.45 3.50 3.70 3.65 3.62 3.32 3.58 3.40 3.43
Mistral 56B 3.88 3.07 2.94 3.30 2.71 3.81 3.75 3.42 2.95 3.94 3.84 3.58 3.31 3.44 3.25 3.33
@@ 4.4.1 CFs for NLI
  • Copy-Paste: When asked to generate a CF, and change the label from contradiction to entailment, LLMs will use the unchanged part (premise or hypothesis) as output. This a clever but lazy way to flip the label to entailment, since two identical sentences would naturally have the label entailment. These CFs were given perfect scores by GPT4. Table 14 in the Appendix shows the percentage of copy-paste CFs for all LLMs (at most 4.27% for GPT3.5).
  • Negation: When asked to to change the label from entailment to contradiction, LLMs would negate the premise/hypothesis. The negation does not make sense in the observed CFs.
  • Inconsistency: These examples contain contradictory or illogical sentences, but GPT4 sometimes incorrectly assigned high scores.
We show examples for each error category in Table 12.
@@ 4.4.2 Evaluation Scores
4.4.3 Bias in GPT4 Scores
Given GPT4’s preference towards its own generations (cf. Table 9), we conduct a qualitative analysis to examine if we agree with the scores given by GPT4 on a set of SA CFs. More specifically, we examine a set of expert CFs that were given lower scores than their corresponding GPT4 CFs on all three metrics. On 12 out of 14 instances we do not agree with the scores given by GPT4. We notice that GPT4 unnecessarily changes some parts of the movie reviews, and introduces changes that do not make sense in the wider context of the reviews. We also noticed that GPT4 changes/omits parts containing potentially harmful content (e.g., torture, sexual content, etc.). Hence, we believe GPT4 prefers its own generations, because these generations do not contain any harmful content (despite safety not being an evaluation criteria).
Original Premise Original Hypothesis Counterfactual Original Label Target Label FL UA RS LLM Changed Part
Copy-Paste
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table on the floor in a building that overlooks the city. contradiction entailment 4.0 4.0 4.0 Mistral 7B hypothesis
Inconsistency
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 Mistral 7B premise
An old man in a blue shirt and white pants glowers by a glass door. A man resting outside A glass man by a door. neutral contradiction 1.0 1.0 1.0 Human Crowd hypothesis
Negation
Two men in costumes with fake carrot noses, top hats, sunglasses and white fur coats that contain white electrical lights. People in costumes Two women in costumes with real carrot noses, no hats, no sunglasses, no coats, no lights. entailment contradiction 1.0 2.0 4.0 LLAMA2 70B premise
Table 12: Categorization of a sample of incorrect NLI CFs with evaluation scores from GPT4.
@@ 5 Related Work
There exists several methods for generating CFs, but most of them are desigend for a specific LLM. The CFs generated by MICE Ross et al. (2021) are intended for debugging models, and not for data augmentation. Polyjuice Wu et al. (2021) requires specifying the type of edits that should be conducted, and the resulting CFs should be manually labeled. Robeer et al. (2021) DISCO Chen et al. (2023) uses GPT-3’s fill-in-the-blanks mode, which is unavailable in most open source LLMs and would require adapting them. CREST Treviso et al. (2023) depends on a rationalizer module and the editor module is a masked LM that needs to be further trained. Instead, we decided to prompt LLMs to generate CFs by providing instructions and an example. We provide more details in Section 3.2.
@@ 6 Conclusion
In this work, we investigated the use of various LLMs for CFs generation. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for SA is less challenging than NLI and HS, where LLMs show weaknesses in generating CFs that change the original label. CFs from LLMs can replace human CFs for the purpose of data augmentation on SA and achieve similar performance, while on NLI and HS further improvements are needed. Further, our results suggest that CFs with minimal changes are essential for data augmentation. We also showed that when asked to asses CFs, LLMs exhibit a strong bias towards agreeing with the provided label even if this label is incorrect. GPT4 appears to be more robust than GPT3.5 against this bias. Furthermore, we showed that GPT4 scores its own generations higher and that safety training might be one reason for this preference, i.e., GPT4 prefers its own generations, because they do not contain any harmful content. Future work should focus on (i) leveraging LLMs for higher quality NLI and HS CFs, which correctly change the label and keep changes minimal, (ii) assessing the evaluation abilities of LLMs in mislabeled data settings, and (iii) investigating the effects of safety training on LLMs as evaluators.
@@ 7 Limitations
Appendix A Successful Generations
Table 13 shows how often LLMs successfully generated CFs, i.e., how often they adhered to the pre-defined template in the prompt.
Test split Success Rate
@@ Appendix A Successful Generations
GPT3.5 100.00
GPT4 99.59
LLAMA2 7B 98.98
LLAMA2 70B 81.76
MISTRAL 7B 84.22
MISTRAL 56B 94.67
@@
GPT3.5 100.00
GPT4 100.00
LLAMA2 7B 96.00
LLAMA2 70B 98.00
MISTRAL 7B 96.12
MISTRAL 56B 99.25
@@
GPT3.5 100.00
GPT4 100.00
LLAMA2 7B 99.62
LLAMA2 70B 94.38
MISTRAL 7B 94.38
MISTRAL 56B 98.25
HS
GPT3.5 63.21
GPT4 76.28
MISTRAL 7B 80.44
MISTRAL 56B 81.44
Table 13: Success rate in generating CFs. We consider generations that do not adhere to the pre-defined template in the prompt as failed generations.
@@ Appendix B Further Analysis
Table 14 shows the percentage of copy/paste examples in NLI CFs. 0
@@
LLAMA2 7B premise 3.00
LLAMA2 7B hypothesis 2.26
LLAMA2 70B premise 2.04
LLAMA2 70B hypothesis 1.06
MISTRAL 7B premise 0.91
MISTRAL 7B hypothesis 1.59
MISTRAL 56B premise 0.63
MISTRAL 56B hypothesis 1.40
@@ Appendix D Examples
Premise Hypothesis Counterfactual Original Label Target Label FL UA RS LLM Changed Part
@@
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 Mistral 7B revised_premise
@@
Appendix E Prompt